OpenStudy (anonymous):

Find the volume of the prism. 192 cu. ft. 64√(13) cu. ft. 96 cu. ft.

OpenStudy (anonymous):

are those dimensions, or multiple choice answers?

OpenStudy (anonymous):

multi choice

OpenStudy (mathstudent55):

\(\Large V_{prism} = Bh\) where B = the area of the base, and h = height of the prism

OpenStudy (mathstudent55):

You have a triangular prism. The base is a triangle. To find the volume of the prism, you need to find the area of the triangular base and multiply it by the height of the prism.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so From what i can see of it, it is 28 but I'm most likely wrong

OpenStudy (mathstudent55):

It's not 28. We'll do it together. Look at the triangular base. What is the area of the triangle below? 

OpenStudy (anonymous):

13?

OpenStudy (mathstudent55):

\(\Large A_{triangle} = \dfrac{bh}{2} \) How did you get 13? What operations did you do?

OpenStudy (mathstudent55):

That's why we are working together, so you'll understand it.

OpenStudy (anonymous):

operations? Guess I really missed a lot on this subject.....)

OpenStudy (mathstudent55):

The area of a triangle is the base times the height divided by 2. In your triangle, the base is 4 ft, and the height is 6 ft. What is 4 * 6 = ?

OpenStudy (anonymous):

24

OpenStudy (mathstudent55):

Operations are addition, subtraction, multiplication, division. Somehow you arrived at 13. I was just wondering how you got that number. Yes, 24 is correct. That is the base times the height. Now we divide that by 2. What is 24/2 = ?

OpenStudy (anonymous):

I added lol, i got 12

OpenStudy (mathstudent55):

Ok. 12 ft^2 is correct. The area of the base is 12 square feet.

OpenStudy (mathstudent55):

Now we have the area of the base, 12 ft^2. Now we use that number in the formula for the volume. The volume of a prism is the area of the base times the height of the prism. The area of the base is 12 ft=^2. The height of the prism is 8 ft. What is 12 ft^2 * 8 ft ?

OpenStudy (anonymous):

umm 1152??

OpenStudy (mathstudent55):

How did you get that? All you need to do is multiply 12 * 8 Then ft^2 * ft = ft^3, which are proper cubic units for a volume. Just multiply 12 * 8

OpenStudy (anonymous):

ohh i thought 12 was squared

OpenStudy (anonymous):

Im REALLY bad at this

OpenStudy (mathstudent55):

No. The units ft^2 are squared because 12 ft^2 is an area. Remember, we found the area of the base to be 12 ft^2.

OpenStudy (anonymous):

96

OpenStudy (anonymous):

so what do I do now with 96?

OpenStudy (mathstudent55):

Correct. Since it's a volume, and we are dealing with ft, the answer is V = 96 ft^3

OpenStudy (mathstudent55):

You can also write 96 ft^3 as 96 cu ft. It just means 96 cubic feet.

OpenStudy (anonymous):

okay thanks.. One more ?

OpenStudy (mathstudent55):

You're welcome. Sure, but pls start a new post.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

NM I have to later not working right now...